# AMEX: Android Multi-annotation Expo Dataset for Mobile GUI Agents

### v1

@@ Abstract
AI agents have drawn increasing attention mostly on their ability to perceive environments, understand tasks, and autonomously achieve goals. To advance research on AI agents in mobile scenarios, we introduce the Android Multi-annotation EXpo (AMEX), a comprehensive, large-scale dataset designed for generalist mobile GUI-control agents. Their capabilities of completing complex tasks by directly interacting with the graphical user interface (GUI) on mobile devices are trained and evaluated with the proposed dataset. AMEX comprises over 104K high-resolution screenshots from 110 popular mobile applications, which are annotated at multiple levels. Unlike existing mobile device-control datasets, e.g., MoTIF [2], AitW [19], etc., AMEX includes three levels of annotations: GUI interactive element grounding, GUI screen and element functionality descriptions, and complex natural language instructions, each averaging 13 steps with stepwise GUI-action chains. We develop this dataset from a more instructive and detailed perspective, complementing the general settings of existing datasets. Additionally, we develop a baseline model SPHINX Agent and compare its performance across state-of-the-art agents trained on other datasets. To facilitate further research, we open-source our dataset, models, and relevant evaluation tools. The project is available at https://yuxiangchai.github.io/AMEX/.
@@ 1 Introduction
AI assistants on mobile devices, such as Siri on iPhones, Bixby on Samsung devices, and Xiao AI on Xiaomi smartphones, have become increasingly prevalent in recent years. While these assistants are adept at managing various routine tasks like setting alarms, conducting web searches, and reporting weather conditions, their capabilities are predominantly confined to interacting with system-built applications. Furthermore, although many can interface with third-party apps via APIs, the lack of universal API support across different mobile operating systems often hampers their performance.
In contrast, human users can complete tasks on mobile devices purely based on visual information from the screen. Inspired by this, researchers are exploring alternative approaches [3, 10, 30] that process vision and natural language inputs, specifically screenshots in mobile environments. We refer to these as Mobile GUI-Control Agents, or GUI Agents for short. These agents are designed to manipulate the user interface elements on the screen directly, with the abilities to interpret natural language commands and analyze screenshot layouts and element functionalities, which theoretically enable agents to execute any task on any app.
@@
Table 1: Comparison of AMEX to other datasets. Scale: the number of unique instructions on general third-party apps, average steps per instruction and screenshots. Diversity: screenshot description, element labels, element functionality and the action details for stepwise operation.
Dataset # Unique
General Inst. # Avg
Steps # Screen-
shots Screen
@@
Func. Action
Detail
PixelHelp 187 4.2 $\sim$800
UGIF 480 6.3 $\sim$3.3K
MoTIF 480 4.5 $\sim$21K
AitW 1539 6.5 $\sim$510K *
AitZ 2504 7.5 $\sim$18K * $\blacksquare$
AMEX 2946 12.8 $\sim$104K
@@
However, the latest GUI agents, such as CogAgent [10] and AppAgent [28], exhibit unsatisfactory performance when handling real-world tasks. The primary issues with these agents are their lack of awareness of page layouts and their limited understanding of the functionalities of various user interface elements. These shortcomings are largely due to the absence of a comprehensive, detailed, large-scale dataset. Recent research has introduced several related datasets for the Android operating system, featuring different types of annotations. For instance, AitW [19] is a large-scale dataset annotated with instructional operations and screenshot icon detection. However, the data quality, especially in element-level annotation which is constructed by a pretrained IconNet, is subpar due to the unregulated annotation style. Besides, only 2.8% of instructions are for general third-party apps, not representative of general usage. Additionally, AitW suffers from heavy data redundancy as 94.9% instructions, including both multi-step and single-step, are in the category of “Web-shopping”. To address these issues, AitZ [32] selects a smaller subset of AitW and re-annotates it with screenshot descriptions and Chain-of-Action-Thought (CoAT) annotations. Though the CoAT paradigm is of some help, the dataset scale shrinks substantially to total 18K screen-action annotation pairs. Other datasets [2, 14, 23] also have various limitations, including the number and practicability of instructions, diversity, accuracy and reliability of annotations, coverage of apps and Android versions. Furthermore, these datasets primarily offer supplemental tree-based representations, such as the View Hierarchy. Notably, the View Hierarchy is not universally available across all apps, nor is it readily accessible to average mobile users. This limitation restricts the practical usability and generalizability of the data, as it does not reflect the typical interaction scenarios encountered by most users.
When human users receive an instruction, they perform several steps beyond merely interpreting the instruction. A human user would comprehend the overall page layout and analyze the current page to identify interactive elements, determining which elements can be tapped and where to scroll. The functionality of those elements are then assessed, particularly the clickable ones. For example, a bell icon likely indicates navigation to a notification page, while a magnifier icon might signify a search page or activate a search bar. Equipped with this information and contextual knowledge, the user can effectively break down the instruction into specific actions for multiple screens.
To align an GUI agent with humans, we design a comprehensive multi-level annotation dataset for the Android OS, named the Android Multi-annotation EXpo (AMEX). AMEX consists of three levels of annotations (see Figure 1): (i) GUI interactive element grounding, (ii) GUI screen and element functionality descriptions, and (iii) instructions with GUI-action chains. In total, AMEX consists of annotations on over 104K high-resolution screenshots, 711K element-wise functionalities, and around 3,000 unique instructions with stepwise GUI-action chains, with an average of 13 steps per instruction (see Table 1 and Table 2). The bounding boxes of GUI interactive elements are filtered and verified by human annotators. Screen descriptions and element functionalities, generated by GPT with built-in app descriptions, are also manually checked. Given random apps and instructions, annotators, all trained with specific and precise guidelines for each action, conduct the operations in a natural manner to achieve the objective. The data is collected from 110 third-party apps with Android 13 on Pixel 7 Pro and Samsung S10 emulators, ensuring up-to-date and prevalent screen resolutions and aspect ratios. The key features of AMEX can be summarized as providing knowledge of the modern mobile GUI environment from multiple levels, and complex human logics and operations on third-party apps.
Our contributions can be summarized as follows: (a) We collect and release AMEX for training and evaluating generalist mobile GUI agents, which has multi-level annotations, providing reliable understandings and complex instructions of the smartphone UI environment; (b) We release the SOTA agent SPHINX Agent, which can serve as the baseline model for future researches on GUI agents.
2 Related Work
@@
GUI-Control Datasets. Table 1 compares several popular GUI-control datasets on Android. While some works [5, 15, 20] focus on the web platform, on Android OS, many works [19, 29, 33] have focused on identifying various types of GUI elements, often assigning numerous classes to different elements. Other studies [2, 14, 23] primarily emphasize action-observation pairs during instructional operations, but their annotations are limited and often require supplemental View Hierarchy (VH) data for each screenshot. Furthermore, these instructions are typically too simplistic for real-world tasks. AitW [19] provides both screen GUI element annotations and instructions, but it includes only a small portion of instructions for third-party apps, with most operations conducted on Chrome and other system-built apps. Additionally, each instruction is repeated multiple times, resulting in significant data redundancy. It also relies on the pre-trained IconNet for automatic annotations, which unfortunately leads to numerous mis-annotated types and misaligned bounding boxes. In response, AitZ [32] filters AitW thoroughly, selecting 2.5K unique instructions and episodes, and introduces the Chain-of-Action-Thought framework to annotate action results and page descriptions better. Despite this refinement, AitZ contains only 18K screen-action pairs.
@@
GUI-Control Agents. Recent advancements have leveraged the extensive world knowledge and robust embodied capabilities of Large Language Models (LLMs) [1, 8, 9, 11, 12, 22] for complex task planning and reasoning within GUIs. A notable approach involves employing business-level generalist models like GPT-4v directly as GUI-control agents. Works [28, 34] have utilized these models, employing extensive prompt engineering to guide the LLM in executing complex tasks. However, the effectiveness of these methods is inherently limited by the capabilities of the available generalist MLLMs [4, 13, 16, 24] and requires meticulous prompt design to achieve optimal results. Alternatively, another research line focuses on fine-tuning smaller LLMs on GUI-specific datasets to imbue them with domain-specific knowledge, thereby enhancing their operational efficiency. For example, CogAgent [10] enhances performance in GUI-related tasks by integrating a high-resolution cross-module that fuses image features from various levels. Similarly, MobileAgent [6] improves sample handling and structuring of input data to make it more consistent and LLM-friendly. Auto-UI [30] utilizes a multimodal encoder-decoder language model based on T5 and BLIP and CoCo-Agent [17] takes element OCR layouts as additional input to enhance the performance. Expanding beyond traditional mobile phone GUI domains, recent studies [26, 31] are exploring the potential of LLM-Agents in operating system-level task execution.
3 Android Multi-annotation EXpo (AMEX)
@@
Figure 3: Overview of the data collection pipeline. The raw data is from two subsets collected by human annotators and an autonomous tool respectively, while annotators record the GUI-action chains simultaneously. Then raw data is sent to annotators to filter GUI element bounding boxes, then they with raw screenshots are sent as input to GPT to extract the GUI screen and element descriptions, which are then manually checked by humans.
When receiving an instruction, a human user first analyzes the overall screen layout to form a basic understanding of the current Android environment. The user then identifies interactive elements and areas, and assesses the functionalities of those elements. Finally, the user breaks down the instruction into simple, step-by-step actions on each screen. Based on this human cognitive process, we design three levels of annotations in our proposed AMEX training set: (i) GUI interactive element grounding (see Section 3.2), (ii) GUI screen and element descriptions (see Section 3.3), and (iii) instructions with GUI-action chains (see Section 3.4).
The raw data is collected using Android emulators, specifically Android Virtual Device and Genymotion [^1] https://www.genymotion.com/. We developed tools to autonomously perform emulator operations as well as record human actions. These tools leverage Appium [^2] https://appium.io/docs/en/latest/, an open-source, cross-platform test automation tool for Android, iOS, and web apps. Appium collects device screenshots during operations, each of which also corresponds to an Extensible Markup Language (XML) file recording the basic screen layout with element attributes, such as bounding boxes and in-app descriptions.
@@ 3.1 Data Collection Pipeline
An overview of the data collection pipeline is illustrated in Figure 3. The raw data is collected through two methods: human instruction-following GUI manipulations and autonomous GUI controls. Human GUI manipulations involve recording stepwise operations for each instruction, and simultaneously storing screenshots and each screen’s XML data before each stepwise operation. In parallel, an autonomous script controls emulators to collect additional screenshots and their XMLs. These two subsets comprise the entire raw dataset. Then for each screenshot, initial bounding boxes of interactive elements and their in-app descriptions (if available) are parsed from the corresponding XML. Human annotators then review each screenshot to filter out all the misaligned boxes, which serve as the interactive element grounding annotations. With the in-app descriptions, GPT generates the functionalities of the selected elements and provides descriptions for the whole screenshot. Human annotators then further check the descriptions of functionalities. More detailed methods are discussed in the following sections and Appendix A.1.3.
@@ 3.2 Level I: GUI Interactive Element Grounding
Existing datasets [19, 32] typically classify elements on the screen, such as icons, texts, and images, based on their types. Instead of adhering to the traditional classification paradigm, we define interactive elements more broadly as any elements that users can interact with, regardless of their specific types (see Figure 2(a)). Specifically, interactive elements in our dataset are only categorized into two subsets: (i) clickable elements and (ii) scrollable elements.
Clickable Elements are the most common components in a screen. They typically include clickable icons, images, texts, and compounds that combine several categories. Figure 2(b) illustrates various clickable elements in red boxes. We also include certain “typeable” elements, such as search bars, because most typeable elements require a prior click action to enable typing.
Scrollable Elements typically occupy larger areas on the screen. In most cases, a scrollable element area supports a pair of actions, such as “scroll down” and “scroll up” or “scroll left” and “scroll right.” Generally, the vertical scrollable area forms the main frame in the middle of the screen, while the horizontal scrollable area serves as a “carousel” for listing different categories or activities. Figure 2(b) illustrates these two types of scrollable elements.
As mentioned in Section 3.1, the collection of screenshots and their XML data utilizes two methods. Here, we provide a more detailed pipeline for executing an autonomous script to perform operations on an Android emulator. The script is designed to traverse an app in an unconstrained manner and collect data. It performs actions (see Appendix A.1.1) at regular time intervals to allow each page to fully load. After waiting, the script captures the current screenshot along with the corresponding XML file. To ensure efficiency and prevent redundancy, we limit the maximum number of operation steps per script execution. Additionally, a break mechanism is implemented to exit potential dead loops, ensuring the script progresses effectively. Upon the raw data (yellow part in Figure 3), we extract bounding boxes for interactive elements from the XML associated with each screenshot. While scrollable elements are typically parsed with high quality, clickable elements are more problematic. This issue is from the fact that app developers often design elements with overlapping or layered coverage, leading to the generation of bounding boxes even for elements that are visually hidden from the user’s view. To address this, annotators manually inspect each screenshot, identifying and selecting only the interactive elements that are actually visible within the interface (see Appendix A.1.2).
@@ 3.3 Level II: GUI Screen and Element Functionality Descriptions
Previous works on GUI elements often rely on predefined class names to convey the underlying meaning of each element. However, this classification-based method has significant limitations. For instance, a small triangle symbol typically indicates “play” in video or music interfaces, but it might also appear as part of a content creator’s profile image. In such cases, the model erroneously interprets the symbol and repeatedly attempts to tap the icon to play the video. This issue arises because the class-based annotation approach focuses on class labels rather than truly understanding the functionality of each element in the surrounding context.
To ensure the dataset is truly instructive rather than merely icon detection, we focus on describing screen status and element functionalities. Consider the above mentioned example: instead of providing a bounding box for the small triangle and labeling it as “ICON_PLAY,” we present the actual functionality of the element within its context (e.g., “Opens the creator’s personal page”). This strategy offers clear, instructive, and detailed description, enhancing the dataset’s applicability. We adopt GPT-4o for the element-wise annotation to describe screens and interpret element functionalities. Given the compact and dense nature of GUI elements, we apply Set of Markers (SoM) techniques [27] to bolster GPT-4o’s capability for visual localization. The coordinates of these elements are obtained and cleaned in the preceding phases (Section 3.2). Additionally, recognizing that some elements possess abstract functionalities that are challenging to discern, we supplement the prompts for GPT-4o with in-app descriptions (if available) extracted from XMLs to enhance GPT-4o’s comprehension of the screen. After the data generation, human annotators manually check 20% of the descriptions of the element functionality. Further elaboration on this methodology is provided in Appendix A.1.3.

### v2

@@ Abstract
AI agents have drawn increasing attention mostly on their ability to perceive environments, understand tasks, and autonomously achieve goals. To advance research on AI agents in mobile scenarios, we introduce the Android Multi-annotation EXpo (AMEX), a comprehensive, large-scale dataset designed for generalist mobile GUI-control agents which are capable of completing tasks by directly interacting with the graphical user interface (GUI) on mobile devices. AMEX comprises over 104K high-resolution screenshots from popular mobile applications, which are annotated at multiple levels. Unlike existing GUI-related datasets, e.g., Rico Deka et al. (2017), AitW Rawles et al. (2024b), etc., AMEX includes three levels of annotations: GUI interactive element grounding, GUI screen and element functionality descriptions, and complex natural language instructions with stepwise GUI-action chains. We develop this dataset from a more instructive and detailed perspective, complementing the general settings of existing datasets. Additionally, we finetune a baseline model SPHINX Agent and illustrate the effectiveness of AMEX.
AMEX: Android Multi-annotation Expo Dataset for Mobile GUI Agents
Yuxiang Chai1*, Siyuan Huang2*, Yazhe Niu1, Han Xiao1, Liang Liu3, Guozhi Wang3, Dingyu Zhang1, Shuai Ren 3, Hongsheng Li 1, 1MMLab @ CUHK, 2SJTU, 3vivo AI Lab, *Equal contribution, Corresponding author: hsli@ee.cuhk.edu.hk
@@ 1 Introduction
In recent years, AI-powered virtual assistants, such as Siri, Bixby, and Xiao AI, have evolved as key tools for facilitating interactions between users and mobile devices. These assistants have demonstrated to be effective in managing routine tasks such as setting alarms, performing web searches, and reporting weather conditions. However, their functionality is often restricted to system-built applications or third-party apps supported by application programming interfaces (APIs). This limitation highlights a significant gap between the capabilities of current AI assistants and the diverse, open-ended ways humans interact with mobile devices.
Unlike AI assistants, human users can accomplish tasks on mobile devices by relying solely on visual information from the screen. Humans intuitively interpret graphical user interfaces (GUIs), analyze page layouts, and infer the functionalities of interactive elements to complete tasks in various apps. Inspired by this human ability, researchers are exploring new paradigms for mobile interaction that leverage vision and natural language processing. One such paradigm involves the development of Mobile GUI-Control Agents, or GUI Agents, which aim to directly interact with screen elements based on visual and textual inputs, such as screenshots and natural language instructions. These agents hold the potential to transcend the limitations of traditional API-based assistants, enabling more flexible and universal interactions with any app or mobile interface.
Existing GUI agents show promise but face significant challenges in real-world applications. Their effectiveness is possibly limited by a lack of understanding of GUI layouts and the functionalities of interactive elements through the observation of wrong predictions. These limitations largely arise from the absence of comprehensive datasets that adequately reflect the complexity and diversity of mobile GUI environments. While several instructional datasets Burns et al. (2021); Sunkara et al. (2022); Gubbi Venkatesh et al. (2024); Rawles et al. (2024b) have been introduced to address this issue, they struggle with limitations such as inaccurate annotations, insufficient task diversity, and a lack of representativeness for general mobile usage. Similarly, other efforts Deka et al. (2017); Li et al. (2020); Feng et al. (2024); Bunian et al. (2021) that focus on GUI element annotation are constrained by limited dataset scale, outdated app versions, and traditional annotation styles that do not provide sufficient information for GUI agents.
@@
Human users approach GUI-based tasks by integrating multiple cognitive processes. They interpret the overall layout of the page, identify interactive elements and assess their functionality, such as recognizing that a bell icon likely leads to notifications and a magnifier icon likely triggers a search function. Users then decompose natural language instructions into actionable steps, navigating across multiple screens to achieve their goals. To design effective GUI agents, it is essential to replicate these human cognitive processes, which requires datasets that comprehensively capture the structure, functionality, and interaction patterns of mobile GUIs.
To address this need, we introduce a new dataset, the Android Multi-annotation EXpo (AMEX), specifically designed to advance the development of GUI agents by providing a multi-level understanding of mobile GUIs. AMEX includes three levels of annotations: (i) GUI interactive element grounding, (ii) GUI screen and element functionality descriptions, and (iii) instructions with GUI-action chains. The dataset comprises over 104K high-resolution screenshots, 21K screen descriptions with 300K element-wise functionalities, and approximately 3,000 unique complex instructions, with an average of 12.8 steps (see Figure 1, Table 1, and Table 2). To ensure annotation precision and quality, all annotations, including bounding boxes, screen and element functionalities, and instructions with GUI-action chains, are verified by human annotators trained with detailed guidelines.
Our contributions can be summarized as follows: (a) We collect and release AMEX, a multi-level dataset, providing reliable and detailed understandings of the smartphone GUI environment, which can serve as a strong supplementary dataset to boost the performance of GUI agents; (b) We train SPHINX Agents, which can serve as the baseline models for future researches on GUI agents and illustrates the effectiveness of AMEX.
2 Related Work
2.1 GUI-related Datasets
Dataset Screen
@@ 2.1 GUI-related Datasets
Func. Task &
Action # Screen-
shots # Element
Func. # Unique
General Inst. # Avg
Steps
RICO 72K - - -
RICO semantics 72K - - -
VINS 4K - - -
MUD 18K - - -
PixelHelp 800 - 187 4.2
UGIF 3.3K - 480 6.3
MoTIF 21K - 480 4.5
AitW * 510K - 1539 6.5
AitZ * $\blacksquare$ 18K 18K 2504 7.5
AndroidControl * 99K - 15,283 4.8
AMEX 104K 296K 3046 12.8
@@
Table 1: Comparison of mobile GUI-related datasets. Scale: the number of unique instructions on general third-party apps, average steps per instruction, number of screenshots and element functionality descriptions. Diversity: screenshot description, element labels, element functionality and the action details for stepwise operation.
Table 2: AMEX statistics
# Screenshots # Apps # Interactive Elements # Functionalities # Instructions # Avg. Steps
104,876 192 1,659,647 296,075 3046 12.8
Table 1 compares several popular GUI-related datasets on Android. While some works Deng et al. (2024); Liu et al. (2018); Shi et al. (2017); Wu et al. (2023) focus on the web or desktop platforms, on Android OS, many works Deka et al. (2017); Li et al. (2020); Feng et al. (2024); Bunian et al. (2021) have focused on identifying various types of GUI elements, assigning classes to different elements. This annotation style can be defined as a traditional icon classification and detection problem. In many cases, however, the functionality of an icon depends on the context of the GUI and many elements on mobile devices are not an icon and their functionalities are also critical. Besides, most of the datasets Deka et al. (2017); Bunian et al. (2021); Li et al. (2020) are outdated, where the screenshot interfaces are completely different from modern apps. The most recent MUD Feng et al. (2024) only provides 18K screenshots, which are insufficient for understanding the layouts and elements for LLMs.
@@
Other studies Burns et al. (2021); Rawles et al. (2024b); Zhang et al. (2024); Gubbi Venkatesh et al. (2024) primarily emphasize action-observation pairs during instructional operations, but their annotations are limited and often require supplemental View Hierarchy (VH) data for each screenshot. AitW Rawles et al. (2024b) provides both screen GUI element annotations and instructions, but it includes only a small portion of instructions for third-party apps, with most operations conducted on Chrome and other system-built apps. Additionally, each instruction is repeated multiple times, resulting in significant data redundancy. In response, AitZ Zhang et al. (2024) filters AitW thoroughly, selecting 2.5K unique instructions and episodes, and introduces the Chain-of-Action-Thought framework to annotate action results and page descriptions better. Despite this refinement, AitZ contains only 18K screen-action pairs. AndroidControl Li et al. (2024) is another large-scale instruction-based dataset, however, the element-wise annotations contain heavy redundancy and misaligned cases due to the lack of human verification.
Figure 2: Overview of the data collection pipeline. The raw data is from two subsets collected by human annotators and an autonomous tool. Annotators record the GUI-action chains simultaneously while collecting the screenshots and corresponding XMLs. Then raw data is sent to annotators to filter GUI element bounding boxes, and then the boxes with corresponding raw screenshots are sent as input to GPT and Gemini to extract the GUI screen and element descriptions, which are then manually checked by humans.
2.2 GUI-related Systems and Agents
@@
Wang et al. (2023) first starts to apply Large Language Model (LLM) on GUI, but it remains on tasks only interacted on single page, which are more like question-answer tasks rather than end-to-end instructional tasks. Recent advancements have leveraged the extensive world knowledge and robust embodied capabilities of LLMs Gu and Dao (2024); Gur et al. (2023); Touvron et al. (2023) for task planning and reasoning. Works Zhang et al. (2023); Zheng et al. (2024) have utilized business-level models (e.g., GPT-4V), employing extensive prompt engineering to guide the LLM in executing complex tasks. The effectiveness of these methods requires a meticulous prompt design to achieve optimal results. Alternatively, another research line focuses on fine-tuning smaller LLMs on GUI-specific datasets to imbue them with domain-specific knowledge. For example, CogAgent Hong et al. (2024) enhances performance in GUI-related tasks by integrating a high-resolution cross-module that fuses image features from various levels. CoCo-Agent Ma et al. (2024) and AndroidControl Li et al. (2024), unlike other agents taking only screenshots as input, use element layouts from accessibility trees or view hierarchy as additional input to enhance the performance. However, still many apps don’t support accessibility information or only provide very little of that.
@@
3 Android Multi-annotation EXpo (AMEX)
When receiving an instruction, a human user first analyzes the overall screen layout to form a basic understanding of the current Android environment. The user then identifies interactive elements and areas, and assesses the functionalities of those elements. Finally, the user breaks down the instruction into simple, step-by-step actions on each screen. Based on this human cognitive process, we design three levels of annotations in our proposed AMEX training set: (i) GUI interactive element grounding (see Section 3.2), (ii) GUI screen and element descriptions (see Section 3.3), and (iii) instructions with GUI-action chains (see Section 3.4). The overall statistics of AMEX are listed in Table 2.
@@ 3.1 Data Collection Pipeline
An overview of the data collection pipeline is illustrated in Figure 2. The raw data is collected through two methods: human instruction-following GUI manipulations and autonomous GUI controls. Human GUI manipulations involve recording stepwise operations for each instruction, and simultaneously storing screenshots and each screen’s Extensible Markup Language (XML) data. In parallel, an autonomous script controls emulators to collect additional screenshots and their XMLs. These two subsets comprise the entire raw dataset. Then for each screenshot, initial bounding boxes of interactive elements and their in-app descriptions (if available) are parsed from the corresponding XML. Human annotators then review each screenshot to filter out all the misaligned boxes, which serve as the interactive element grounding annotations. With the in-app descriptions, GPT and Gemini generates the functionalities of the selected elements and provides descriptions for the whole screenshot. Annotators then further check the quality of the descriptions of functionalities. More details are discussed in the following sections and Appendix A.1.
@@ 3.2 Level I: GUI Interactive Element Grounding
Existing datasets Deka et al. (2017); Sunkara et al. (2022); Feng et al. (2024); Bunian et al. (2021) typically classify elements on the screen, such as icons, texts and images, based on their types. Instead of adhering to the traditional classification paradigm, we define interactive elements more broadly as any elements that users can interact with, regardless of their specific types (see Figure 3(a)). Specifically, interactive elements in our dataset are only categorized into two subsets: (i) clickable elements and (ii) scrollable elements. The motivation of our annotation is mainly from two facts: (i) no existing dataset provides information on the “scroll” action, which is crucial especially for agents to know the horizontally scrollable area. (ii) “click” action is the most used action and many “compound” elements cannot be simply categorized. As illustrated in Figure 3(b) red boxes, four “compound” elements in the middle contain both image and text and the definition of “clickable” is more suitable with the functionality in Level 2 for agents to understand the element.
Clickable Elements are the most common components in a screen. They typically include clickable icons, images, texts, and compounds that combine several categories. We also include certain “typeable” elements, such as search bars, because most typeable elements require a prior click action to enable typing.
Scrollable Elements typically occupy larger areas on the screen. In most cases, a scrollable element area contains many clickable elements and supports a pair of actions, such as “scroll down” and “scroll up” or “scroll left” and “scroll right.” Figure 3(b) illustrates these two types of scrollable elements.
@@ 3.3 Level II: GUI Screen and Element Functionality Descriptions
Previous works Deka et al. (2017); Feng et al. (2024); Bunian et al. (2021) on GUI elements often rely on predefined class names, such as “image” and “text”, to convey the underlying meaning of each element. Rico Semantics Sunkara et al. (2022) annotates the icons with slightly more detailed semantics information but the annotations are still classifications on icons. However, this classification-based method has significant limitations. For instance, a “plus” symbol typically indicates “plus” in calculator interfaces, but it means “create a new task” in a ToDo interface. This class-based annotation approach focuses on class labels rather than truly understanding the functionality of each element in the surrounding context, which may lead to a misunderstanding of the page layout.
To ensure the dataset is truly instructive rather than merely icon detection, we focus on describing screen status and element functionalities. Consider the above mentioned example: instead of providing a bounding box for the icon and labeling it as “ICON_PLUS”, we present the actual functionality of the element within its context (e.g., “Create a new task”). This strategy offers a clear, instructive, and detailed description of the interactive element, enhancing the dataset’s applicability. Besides, the traditional categories of “IMAGE” and “TEXT” provides very little information of the actual element functionality. Our annotation will provide a much more detailed and informative functionality for those elements (e.g., “View the detail page of the ’Jazz Classics’ playlist” for the red left bottom “compound” element in Figure 3(b)). Further elaboration on the collection and processing is provided in Appendix A.1.4.